OpenStudy (anonymous):

I don't understand, is there a certain formula I have to use. Find the surface area of the square pyramid. Round your answer to the nearest hundredth

OpenStudy (mathmale):

Dear Laughs, Your square pyramid has four triangular surfaces. You'll need to find the area of one such surface, and then multiply that area by four.

OpenStudy (anonymous):

The formula to find the surface area of a square pyramid is: \[A = a^{2} + 2a \sqrt{\frac{ a^{2} }{ 4 } + h}\]

OpenStudy (anonymous):

I think that's right... it's been a while since I've done surface area. Do you know if it is mathmale?

OpenStudy (mathmale):

Do you have a diagram showing the pyramid? The area of a triangle with base B and height H is (1/2)B*H. I've never used the formula \[A = a^{2} + 2a \sqrt{\frac{ a^{2} }{ 4 } + h}\]myself, but see that that a^2 is the area of the base of the pyramid.

OpenStudy (mathmale):

In your shoes, I'd either figure out the surface area of one triangular side of this pyramid, multiply that by four, and then add the area of the base (a^2)...or, I'd do an Internet search for "surface area of a square pyramid."

OpenStudy (anonymous):

That's what I was taught in my math book so I assume it's correct, or at least I hope it is or else i've learned SA of a square pyramid wrong lol

OpenStudy (mathmale):

I'm sure we could find that formula, or derive it...but if we're not going to be finding the surface area of square pyramids often, I'd vote for this "find the surface area of one triangle" approach.

OpenStudy (anonymous):

The method you mentioned first is probably an easier route to go, or at least it makes sense to me.

OpenStudy (mathmale):

@AnyTipzWouldHelp 's memory is working very well. I've verified the formula by doing an internet search: https://www.google.com/search?q=surface+area+of+a+square+pyramid&oq=surface+area+of+a+square+pyramid&aqs=chrome..69i57j0l5.11547j0j7&sourceid=chrome&es_sm=0&ie=UTF-8

OpenStudy (mathmale):

So, Laughs, you pays your money and you takes your choice (as the saying goes).

OpenStudy (anonymous):

Well, it was just last year that I learned it so it wasn't that far away for me to remember :P

OpenStudy (mathmale):

Pretty cool. And here I am, a retired math teacher, and don't even know that formula. :(

OpenStudy (mathmale):

Laughs??

OpenStudy (mathmale):

To use either method, you'll need to know the length of one side of the square base and the height of the pyramid. Do you have that info?